Ten Pieces of paper numbered consecutively from 1 to 10 are put into a hat. If three pieces of paper are drawn at random from the hat one at a time without being returned to the hat, What is the probability that all three will be Even numbers
Answer by stanbon(75887)

Ten Pieces of paper numbered consecutively from 1 to 10 are put into a hat. If three pieces of paper are drawn at random from the hat one at a time without being returned to the hat, What is the probability that all three will be Even numbers
-------------
5 of the numbers are even.
---
# of ways to select 3 even is 5C3 = 5C2 = 10
---
# of ways to select 3 randomly 10C3 = 128
============
P(3 even) = 10/128 = 5/63
===============================
Cheers,
Stan H.
============